Write z1=3–√+i and z2=1−i in trigonometric form. Then the argument of the quotient z1/z2is given by

1
Expert's answer
2022-06-13T05:12:03-0400
"r^2=(-\sqrt{3})^2+(1)^2=4""r\ge0=>r=\sqrt{4}=2"

Quadrant II


"\tan \theta=\dfrac{-1}{\sqrt{3}}=-\dfrac{1}{\sqrt{3}}""\theta=\pi+\tan^{-1}(-\dfrac{1}{\sqrt{3}})=\dfrac{5\pi}{6}""z_1=2(\cos\dfrac{5\pi}{6}+i\sin\dfrac{5\pi}{6})"





"r^2=(1)^2+(-1)^2=2""r\ge0=>r=\sqrt{2}"

Quadrant IV


"\tan \theta=\dfrac{-1}{\sqrt{3}}=-\dfrac{1}{1}=-1""\theta=2\pi+\tan^{-1}(-1)=\dfrac{7\pi}{4}""z_2=\sqrt{2}(\cos\dfrac{7\pi}{4}+i\sin\dfrac{7\pi}{4})"

"z_1\/z_2"



"\theta_{z_1/z_2}=\dfrac{5\pi}{6}-\dfrac{7\pi}{4}=-\dfrac{11\pi}{12}"